OpenStudy (anonymous):

the cost of 3 apples and 5 pears was 280 dollars. if the apple's cost increased by 10 dollars and the pears cost stayed the same, after the change; 2 apples and 4pears were 220 dollars. what was the cost of one pear and the cost of one apple? ~sorry for the english.

OpenStudy (anonymous):

that is alot of money for those apples and pears

OpenStudy (anonymous):

its translated xD i changed the currency too ^-^

OpenStudy (ajprincess):

Say that the price of an apple is x and the price of a pear is y. Then 3x+5y=280 Nw after increasing the price of apple by 10 the price of an apple will be x+10 Then 2(x+10)+4y=220 2x+20+4y=220 x+10+2y=110 x+2y=100 Solve the simultaneous eqns x+2y=100 3x+5y=280 to get the price of apple and pear

OpenStudy (anonymous):

thank you!

OpenStudy (ajprincess):

yw.:)